# Asymptotic error distribution for stochastic Runge–Kutta methods of strong order one

### v1

@@ Abstract.
This work gives the asymptotic error distribution of the stochastic Runge–Kutta (SRK) method of strong order $1$ applied to Stratonovich-type stochastic differential equations. For dealing with the implicitness introduced in the diffusion term, we provide a framework to derive the asymptotic error distribution of diffusion-implicit or fully implicit numerical methods, which enables us to construct a fully explicit numerical method sharing the same asymptotic error distribution as the SRK method. Further, we show for the multiplicative noise that the limit distribution $U(T)$ satisfies $\mathbf{E}|U(T)|^{2}\leq e^{L_{1}T}(1+\eta_{1})T^{3}$ for some $\eta_{1}$ only depending on the coefficients of the SRK method. Thus, we infer that $\eta_{1}$ is the key parameter reflecting the growth rate of the mean-square error of the SRK method. Especially, among the SRK methods of strong order $1$, those of weak order $2$ ($\eta_{1}=0$) share the unified asymptotic error distribution and have the smallest mean-square errors after a long time. This property is also found for the case of additive noise. It seems that we are the first to give the asymptotic error distribution of fully implicit numerical methods for SDEs.
@@ 1. Introduction
where $\mathbf{W}=(W_{1},\ldots,W_{m})$ is an $m$-dimensional standard Brownian motion and $f:\mathbb{R}^{d}\to\mathbb{R}^{d},\,g\in\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$ are Lipschitz functions such that (1.1) admits a unique strong solution. Stochastic numerical methods have been playing the important role in approximating the solutions of SDEs and modeling the intrinsic behaviors of SDEs over the past decades, which have received extensive attention and study (cf. [3, 7, 8, 13, 17]). In the theoretical study on stochastic numerical methods, the error analysis, measuring the reliability and accuracy of stochastic numerical methods, is one of the most basic and most important subject. Compared with the fruitful results on the convergence analysis of stochastic numerical methods, including the strong convergence, weak convergence and the convergence in probability, the investigation on probabilistic limit theorems for errors between numerical solutions and exact solutions has been far from sufficient. In this paper, we are interested in the asymptotic error distribution of stochastic numerical methods applied to the SDE (1.1), which is a kind of generalized central limit theorem.
@@
Asymptotic error distribution of stochastic numerical methods is the limit distribution of the normalized error between numerical solutions and exact solutions as the step-size tends to zero, where the normalized error is the error weighted by the corresponding strong convergence order of numerical methods. The asymptotic error distribution provides the optimal strong convergence order for numerical methods and characterizes the evolution pattern of distribution of the error in the small step-size regime. It also finds some applications in the error structure of numerical methods (cf. [1]). The study on the asymptotic error distribution dated back to the work [14] for SDEs with bounded coeffcients, where the authors gave a sufficient condition for Euler method to admit the asymptotic error distribution. For SDEs with linearly growing coefficients, [11] gave a sufficient and necessary condition for Euler method to admit the asymptotic error distribution. Further, [20] established the asymptotic error distribution of Euler method for SDEs with locally Lipschitz coefficients. In addition, the asymptotic error distribution of $\theta$-method is obtained in [6]. We also refer to the asymptotic error distribution of numerical methods applied to stochastic integral equations ([4, 19]), Mckean-Vlasov SDEs ([15]) and SDEs driven by fractional Brownian motions [9, 18, 23].
@@
Most of the aforementioned works consider the asymptotic error distribution of explicit methods for SDEs. To the best of our knowledge, in the existing literature, there are no results on the asymptotic error distribution of diffusion-implicit numerical methods (the implicitness is introduced in the diffusion term $g$) for (1.1). In this work, we are devoted to filling this gap. More precisely, we study the asymptotic error distribution of a class of SRK methods of strong order $1$ applied to (1.1). To present our arguments clearly for dealing with the asymptotic error distribution of diffusion-implicit numerical methods, we will focus on the scalar noise case (i.e., $m=1$) and study the SRK methods (3.9). Generally speaking, when establishing the asymptotic error distribution of a numerical method for (1.1), one usually finds a continuous version of the numerical method and properly decomposes the normalized error into some main terms and remainder terms. Then Jacod’s theory of convergence in distribution of conditional Gaussian martingales (cf. [4, 10, 11]) will play a key role in studying the convergence in distribution of some stochastic integrals in main terms. However, for a diffusion-implicit or fully implicit method, the implicitness introduced in the diffusion term makes it not easy to find a proper continuous version of the numerical method, since one must consider the adaptedness of integrands in stochastic integrals. In addition, diffusion-implicit methods usually introduce the truncated Brownian increments to ensure the solvability, which also makes some trouble for the continuization of numerical methods and dealing with the convergence in distribution of stochastic integrals.
@@
In order to circumvent the difficulty of directly making the continuization of the SRK method (3.9), we are devoted to finding a fully explicit numerical method which shares the same asymptotic error distribution as the aim method. In order to fulfill this, the explicit numerical method needs to approximate the exact solution of (1.1) with strong order $1$ and approximate the numerical solution of the SRK method with strong order no less than $1.5$. We call this explicit numerical method an appurtenant method of the SRK method (3.9). In order to determine the appurtenant method, we adapt the fundamental strong convergence theorem [22, Theorem 2.1] to give the strong error order between two Markov chains (see Theorem 2.1). Based on Theorem 2.1 and the deterministic Taylor expansion, the appurtenant method can be found. Thus, we indeed give a framework to study the asymptotic error distribution of a diffusion-implicit or fully implicit numerical method $\{Y_{n}\}_{n=0}^{N}$, which is summarized as follows.
Framework $1$ for establishing the asymptotic error distribution of $Y_{N}$.
1. Determine the strong convergence order $p$ of $\{Y_{n}\}_{n=0}^{N}$.
@@
2. Find the appurtenant method $\{\widetilde{Y}_{n}\}_{n=0}^{N}$ based on Theorem 2.1 and the Taylor expansion.
Here, $\{\widetilde{Y}_{n}\}_{n=0}^{N}$ is required to approximate the solution to (1.1) with strong order $p$ and
approximates $\{Y_{n}\}_{n=0}^{N}$ with strong order $p+\frac{1}{2}$.
3. Construct a proper continuization of $\{\widetilde{Y}_{n}\}_{n=0}^{N}$ and establish the limit distribution of
$N^{p}(\widetilde{Y}_{N}-Y(T))$, which is the limit distribution of $N^{p}({Y}_{N}-Y(T))$.
@@
In Framework 1, the asymptotic error distribution of ${Y}_{N}$ is the limit distribution of $N^{p}({Y}_{N}-Y(T))$, which equals to $N^{p}({Y}_{N}-\widetilde{Y}_{N})+N^{p}(\widetilde{Y}_{N}-Y(T))$. Note that $N^{p}({Y}_{N}-\widetilde{Y}_{N})$ converges to 00 in probability since ${Y}_{N}-\widetilde{Y}_{N}$ is of strong order $p+\frac{1}{2}$. According, $N^{p}({Y}_{N}-Y(T))$ and $N^{p}(\widetilde{Y}_{N}-Y(T))$ have the same limit distribution due to Slutzky’s theorem (cf. [12, Theorem 13.18]).
@@
  • We provide a framework (Framework 1) to establish the asymptotic error distribution of diffusion-implicit or fully-implicit numerical methods for SDEs. It seems that we are the first to give the asymptotic error distribution of fully implicit numerical methods for (1.1). It should be mentioned that although we only present the results for the SRK method of strong order $1$, Framework 1 applies to higher-order methods.
  • We give the key parameter reflecting the growth rate of the mean-square error of the SRK method by analyzing the properties of the limit distribution. Our results reveal that among the SRK methods of strong order $1$, those of weak order $2$ have the smallest mean-square errors after a long time. Especially, the asymptotic error distribution of the considered SRK methods takes the unified form, when endowed with second order weak convergence.
The rest of the article is organized as follows. We give some notations and Theorem 2.1 for estimating the strong error order between two Markov chains in Section 2. Sections 3 and 4 are devoted to establishing the asymptotic error distribution of the SRK method for (1.1) with multiplicative noise, by using Framework 1. Section 5 gives the asymptotic error distribution of the SRK method for the case of additive noise. Finally numerical experiments are performed to verify our theoretical analyses in Section 6.
@@ 2. Preliminaries
Denote by $\mathbf{C}(\mathbb{R}^{d};\mathbb{R}^{m})$ (resp. $\mathbf{C}^{k}(\mathbb{R}^{d};\mathbb{R}^{m})$) the space consisting of continuous (resp. $k$th continuously differentiable) functions from $\mathbb{R}^{d}$ to $\mathbb{R}^{m}$. Let $\mathbf{C}_{b}^{k}(\mathbb{R}^{d};\mathbb{R}^{m})$, $k\in\mathbb{N}^{+}$, stand for the set of $k$th continuously differentiable functions from $\mathbb{R}^{d}$ to $\mathbb{R}^{m}$ with bounded derivatives up to order $k$. When no confusion occurs, $\mathbf{C}(\mathbb{R}^{d};\mathbb{R}^{m})$ (resp. $\mathbf{C}^{k}_{b}(\mathbb{R}^{d};\mathbb{R}^{m})$) is simply written as $\mathbf{C}(\mathbb{R}^{d})$ (resp. $\mathbf{C}^{k}_{b}(\mathbb{R}^{d})$). For $f\in\mathbf{C}^{k}(\mathbb{R}^{d};\mathbb{R})$, denote by $\mathcal{D}^{k}f(x)(\xi_{1},\ldots,\xi_{k})$ the $k$th order Gǎteaux derivative along the directions $\xi_{1},\ldots,\xi_{k}\,\in\mathbb{R}^{d}$, i.e., $\mathcal{D}^{k}f(x)(\xi_{1},\ldots,\xi_{k})=\sum_{i_{1},\ldots,i_{k}=1}^{d}\,\frac{\partial^{k}f(x)}{\partial x_{i_{1}}\cdots\partial x_{i_{k}}}\,\xi_{1}^{i_{1}}\cdots\xi_{k}^{i_{k}}$. When $k=1$, we write $\mathcal{D}^{1}f(x)(\xi_{1})=\mathcal{D}f(x)(\xi_{1})$ for $f\in\mathbf{C}^{1}(\mathbb{R}^{d};\mathbb{R})$, which equals to $\nabla f(x)^{T}\xi_{1}$, where $\nabla f$ is the gradient of $f$. For $f=(f_{1},\ldots,f_{m})^{\top}\in\mathbf{C}^{k}(\mathbb{R}^{d};\mathbb{R}^{m})$, denote $\mathcal{D}^{k}f(x)(\xi_{1},\ldots,\xi_{k})=(\mathcal{D}^{k}f_{1}(x)(\xi_{1},\ldots,\xi_{k}),\ldots,\mathcal{D}^{k}f_{m}(x)(\xi_{1},\ldots,\xi_{k}))^{\top}$. In this case, $\mathcal{D}f(x)(\xi_{1}):=\mathcal{D}^{1}f(x)(\xi_{1})$, which is nothing but $\nabla f(x)\xi_{1}$ with $\nabla f(x)$ being the Jacobian matrix of $f$.
The Gǎteaux derivative of a matrix-valued function is defined as previously. For $f\in\mathbf{C}^{k}(\mathbb{R}^{d};\mathbb{R})$ the notation $\mathcal{D}^{k}f(x)$ is viewed as a tensor, i.e., a multilinear form defined on $(\mathbb{R}^{d})^{\otimes^{k}}$. And $\|\cdot\|_{\otimes}$ denotes the norm of a tensor. Let $\mathbf{F}$ be set of functions growing at most polynomially, i.e., a (tensor-valued) function $f\in\mathbf{F}$ means that there exist $C>0$ and $\eta>0$ such that for any $x\in\mathbb{R}^{d}$, $|f(x)|\leq C(1+|x|^{\eta})$ or $\|f(x)\|_{\otimes}\leq C(1+|x|^{\eta})$. Throughout this paper, let $K(a_{1},a_{2},...,a_{m})$ denote some generic constant dependent on the parameters $a_{1},a_{2},...,a_{m}$ but independent of the step-size $h$, which may vary for each appearance.
Next, we present a theorem for analyzing the strong error order between two Markov chains. Let $\mathbf{W}(t)=(W_{1}(t),\ldots,W_{m}(t))^{\top}_{t\in[0,T]}$ be an $m$-dimensional standard Brownian motion defined on $(\Omega,\mathcal{F},\mathbf{P})$ with respect to (w.r.t) a filtration $\{\mathcal{F}_{t}\}_{t\in[0,T]}$ with $\{\mathcal{F}_{t}\}_{t\in[0,T]}$ satisfying usual conditions.
Let $\{t_{0}<t_{1}<\cdots<t_{N}=T\}$, $N\in\mathbb{N}^{+}$ be a uniform partition of $[t_{0},T]$, i.e., $t_{k}-t_{k-1}=\frac{T-t_{0}}{N}:=h$ for any $k=1,\ldots,N$. Introduce two one-step mappings $\bar{X}_{t,x}(t+h)$ and $\widetilde{X}_{t,x}(t+h)$ depending on $t,x,h$, which are defined respectively

### v2

@@ Abstract.
This work gives the asymptotic error distribution of the stochastic Runge–Kutta (SRK) method of strong order 111 applied to Stratonovich-type stochastic differential equations. For dealing with the implicitness introduced in the diffusion term, we propose a framework to derive the asymptotic error distribution of diffusion-implicit or fully implicit numerical methods, which enables us to construct a fully explicit numerical method sharing the same asymptotic error distribution as the SRK method. Further, we show that the limit distribution $U(T)$ satisfies $\mathbf{E}|U(T)|^{2}\leq e^{L_{1}T}(1+\eta_{1})T^{3}$ for some $\eta_{1}\geq 0$ only depending on the coefficients of the SRK method. Thus, we infer that $\eta_{1}$ is the key parameter reflecting the growth rate of the mean-square error of the SRK method. Especially, among the SRK methods of strong order 111, those of weak order 222 correspond to $\eta_{1}=0$, sharing the unified asymptotic error distribution, and have the smallest mean-square errors after a long time. This property is also found for the case of additive noise. It seems that we are the first to give the asymptotic error distribution of fully implicit numerical methods for stochastic differential equations.
@@ 1. Introduction
where $\mathbf{W}(t)=(W_{1}(t),\ldots,W_{m}(t))^{\top}_{t\in[0,T]}$ is an $m$-dimensional standard Brownian motion defined on $(\Omega,\mathcal{F},\mathbf{P})$ with respect to (w.r.t.) a filtration $\{\mathcal{F}_{t}\}_{t\in[0,T]}$ with $\{\mathcal{F}_{t}\}_{t\in[0,T]}$ satisfying usual conditions, and $f:\mathbb{R}^{d}\to\mathbb{R}^{d},\,g\in\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$ are Lipschitz continuous such that (1.1) admits a unique strong solution. Stochastic numerical methods have been playing an important role in numerically approximating SDEs and modeling the intrinsic behaviors of SDEs over the past decades, which have received extensive attention and study (cf. [12, 16]). In the theoretical study on stochastic numerical methods, the error analysis, measuring the reliability and accuracy of stochastic numerical methods, is one of the most basic and most important subject. Compared with the fruitful results on the convergence analysis of stochastic numerical methods, including the strong convergence, weak convergence and convergence in probability, the investigation on probabilistic limit theorems for errors between numerical solutions and exact solutions has been far from sufficient. In this paper, we are interested in the asymptotic error distribution of stochastic numerical methods applied to the SDE (1.1), which is a kind of generalized central limit theorem.
1.1. Research background
@@
The asymptotic error distribution of stochastic numerical methods is the limit distribution of the normalized error between numerical solutions and exact solutions as the step-size tends to zero, where the normalized error is the error weighted by the corresponding strong convergence order of numerical methods. The asymptotic error distribution provides the optimal strong convergence order for stochastic numerical methods and characterizes the evolution pattern of distribution of the error in the small step-size regime. It also finds some applications in the error structure of stochastic numerical methods (cf. [1]). The study on the asymptotic error distribution dated back to the work [13] for SDEs with bounded coefficients, where the authors gave a sufficient condition for Euler method to admit the asymptotic error distribution. Subsequent research by [10] provided a sufficient and necessary condition for the asymptotic error distribution of the Euler method when applied to SDEs with linearly growing coefficients. Further, [19] established the asymptotic error distribution of the Euler method for SDEs with locally Lipschitz coefficients. Recently, the asymptotic error distribution of $\theta$-method, applied to SDEs with additive noise, was obtained in [7]. We also refer to the asymptotic error distribution of numerical methods applied to stochastic integral equations ([4, 18]), Mckean–Vlasov SDEs ([14]) and SDEs driven by fractional Brownian motions [8, 17, 22].
@@
Most of the aforementioned works consider the asymptotic error distribution of explicit methods for SDEs. To the best of our knowledge, in the existing literature, there are no results on the asymptotic error distribution of diffusion-implicit numerical methods (the implicitness is introduced via the diffusion term g) for (1.1), despite their superior stability and frequent applications in constructing structure-preserving algorithms. The aim of this paper is to fill this gap. More precisely, we study the asymptotic error distribution of a class of SRK methods of strong order 111 applied to (1.1). Generally speaking, when establishing the asymptotic error distribution of a numerical method for (1.1), one usually finds a continuous version of the numerical method and properly decomposes the normalized error into some dominated terms and remainder terms. Then Jacod’s theory on convergence in distribution of conditional Gaussian martingales (cf. [4, 9, 10]) will play a key role in studying the convergence in distribution of stochastic integrals in dominated terms. However, for a diffusion-implicit or fully implicit method, the implicitness introduced in the diffusion term makes it not easy to find a proper continuous version of the numerical method, since the adaptedness of integrands in stochastic integrals must be taken into account. In addition, diffusion-implicit methods usually introduce the truncated Brownian increments to ensure the solvability, which also makes some trouble for the continuization of numerical methods and dealing with the convergence in distribution of stochastic integrals.
1.2. Framework to derive the asymptotic error distribution
@@
Aiming at presenting our arguments clearly for dealing with the asymptotic error distribution of diffusion-implicit numerical methods, we will focus on the scalar noise case (i.e., $m=1$) and multidimensional additive noise, and study the SRK methods (3.9) and (5.2). We note that our framework remains applicable to multidimensional multiplicative noise case, though at the cost of increased computational effort. In order to circumvent the difficulty of directly making the continuization of the SRK method (3.9), our strategy is to find a fully explicit numerical method which shares the same asymptotic error distribution as the SRK method. Such an explicit numerical method needs to approximate the exact solution of (1.1) with strong order 111 and approximate the numerical solution of the SRK method with strong order strictly greater than 111. We call this explicit numerical method an appurtenant method of the SRK method (3.9). In order to determine the appurtenant method, we generalize the fundamental strong convergence theorem [21, Theorem 2.1] to give the strong error order between any two Markov chains (see Theorem 2.1). Based on Theorem 2.1 and the deterministic Taylor expansion, the appurtenant method can be constructed. Thus, we indeed give a framework to study the asymptotic error distribution of a diffusion-implicit or fully implicit numerical method $\{Y_{n}\}_{n=0}^{N}$, which is summarized as follows.
Table 1. Framework 111 for establishing the asymptotic error distribution of $Y_{N}$.
Step 1. Determine the strong convergence order $p$ of $\{Y_{n}\}_{n=0}^{N}$.
@@
Step 2. Find the explicit appurtenant method $\{\widetilde{Y}_{n}\}_{n=0}^{N}$ based on Theorem 2.1
and Taylor expansion. Here, $\{\widetilde{Y}_{n}\}_{n=0}^{N}$ is required to approximate the solution to
(1.1) with strong order $p$ and approximates $\{Y_{n}\}_{n=0}^{N}$ with strong order $q>p$.
Step 3. Construct a proper continuization of $\{\widetilde{Y}_{n}\}_{n=0}^{N}$ and establish the limit
distribution of $N^{p}(\widetilde{Y}_{N}-Y(T))$, which is that of $N^{p}({Y}_{N}-Y(T))$.
@@
In Framework 1, the asymptotic error distribution of ${Y}_{N}$ is the limit distribution of $N^{p}({Y}_{N}-Y(T))$, which equals to $N^{p}({Y}_{N}-\widetilde{Y}_{N})+N^{p}(\widetilde{Y}_{N}-Y(T))$. Note that $N^{p}({Y}_{N}-\widetilde{Y}_{N})$ converges to 0 in probability since ${Y}_{N}-\widetilde{Y}_{N}$ is of strong order $q>p$. Accordingly, $N^{p}({Y}_{N}-Y(T))$ and $N^{p}(\widetilde{Y}_{N}-Y(T))$ have the same limit distribution due to Slutzky’s theorem (cf. [11, Theorem 13.18]).
1.3. Main results and contributions
@@
  • We propose a practical and effective framework (Framework 1) for establishing the asymptotic error distribution of diffusion-implicit or fully-implicit numerical methods for SDEs. In particular, this framework is successfully applied to deriving the asymptotic error distribution of a class of fully implicit methods for SDEs, which, to the best of our knowledge, is reported here for the first time. It is worth mentioning that Framework 1 is promisingly applicable to higher-order methods for SDEs and implicit numerical methods for other types of stochastic systems.
  • We identify the key parameter reflecting the growth rate of the mean-square error of the SRK method by analyzing the properties of the limit distribution. Our results reveal that among the SRK methods of strong order 111, those of weak order 222 have the smallest mean-square errors after a long time. Especially, the asymptotic error distribution of the considered SRK methods takes the unified form, when endowed with second order weak convergence.
  • We are the first to uncover a novel relationship between the asymptotic error distribution of stochastic numerical methods and their weak convergence order. From the main results of the paper, we conjecture that among stochastic numerical methods with same strong order, those with higher weak convergence order will possess smaller strong error in the long-time simulations.
The rest of the article is organized as follows. We establish Theorem 2.1 for estimating the strong error order between two Markov chains in Section 2 and Theorem 2.3 about stable convergence for a class of SDEs. In addition, we establish the convergence in distribution for a class of SDEs in order to handle the asymptotic error distribution for the multiplicative noise case. Sections 3 and 4 are devoted to establishing the asymptotic error distribution of the SRK method for (1.1) with scalar noise, by using Framework 1. Section 5 gives the asymptotic error distribution of the SRK method for the case of additive noise. Several numerical experiments are performed to verify our theoretical analyses in Section 6. Finally, Section 7 gives the summary of this work.
@@ 2. Preliminaries
In this section, we focus on presenting two theorems on fundamental strong convergence results for numerical method and convergence in distribution for a special class of SDEs.
@@
Denote by $\mathbf{C}(\mathbb{R}^{d})$ (resp. $\mathbf{C}^{k}(\mathbb{R}^{d})$) the space of continuous (resp. $k$th continuously differentiable) functions defined on $\mathbb{R}^{d}$. Let $\mathbf{C}_{b}^{k}(\mathbb{R}^{d})$, $k\in\mathbb{N}^{+}$, stand for the subset of $\mathbf{C}^{k}(\mathbb{R}^{d})$, consisting of functions whose derivatives up to order $k$ are bounded. For a real-valued function $f\in\mathbf{C}^{k}(\mathbb{R}^{d})$, denote by $\mathcal{D}^{k}f(x)(\xi_{1},\ldots,\xi_{k})$ the $k$th order Gǎteaux derivative along the directions $\xi_{1},\ldots,\xi_{k}\,\in\mathbb{R}^{d}$. For an $\mathbb{R}^{m}$-valued function $f=(f_{1},\ldots,f_{m})^{\top}\in\mathbf{C}^{k}(\mathbb{R}^{d})$, denote $\mathcal{D}^{k}f(x)(\xi_{1},\ldots,\xi_{k}):=(\mathcal{D}^{k}f_{1}(x)(\xi_{1},\ldots,\xi_{k}),\ldots,\mathcal{D}^{k}f_{m}(x)(\xi_{1},\ldots,\xi_{k}))^{\top}$. Let $\mathbf{F}$ be the set of functions growing at most polynomially, i.e., a (tensor-valued) function $f\in\mathbf{F}$ means that there exist $C>0$ and $\eta>0$ such that for any $x\in\mathbb{R}^{d}$, $|f(x)|\leq C(1+|x|^{\eta})$ or $\|f(x)\|_{\otimes}\leq C(1+|x|^{\eta})$, where $\otimes$ denotes the norm of a tensor. Throughout this paper, let $K(a_{1},a_{2},...,a_{m})$ be some generic constant dependent on the parameters $a_{1},a_{2},...,a_{m}$ but independent of the step-size $h$, which may vary for each appearance.
2.1. Fundamental strong convergence theorem
In this part, we present a theorem for analyzing the strong error order between two Markov chains, which will be used to construct the appurtenant method of (3.9). Let $\{t_{0}<t_{1}<\cdots<t_{N}=T\}$, $N\in\mathbb{N}^{+}$ be a uniform partition of $[t_{0},T]$, i.e., $t_{k}-t_{k-1}=\frac{T-t_{0}}{N}:=h$ for any $k=1,\ldots,N$. Introduce two one-step mappings $\bar{X}_{t,x}(t+h)$ and $\widetilde{X}_{t,x}(t+h)$ depending on $t,x,h$, which are defined respectively by